OpenStudy (anonymous):

Which of the following fractions is not equivalent 8/22. 4/11 2/5 12/33 16/44

OpenStudy (haseeb96):

it is a clue for u that in 8/22 numenator and denomenator both are even so guess a answer which has different

OpenStudy (jojo4eva):

The answer would be 4/11

OpenStudy (holly00d1248):

No thats not the answer 4/11 is equivilant

OpenStudy (anonymous):

4/11 and 16/44 are easy because one is multiplied simply by two and the other is divided by 2

OpenStudy (jojo4eva):

4/11 is the reduced version of 8/22. Divide the numerator and denominator by 2 and you get 8/2 over 22/2 and it equals 4/11

OpenStudy (jojo4eva):

oh it would be 2/5 and 12/33 then

OpenStudy (holly00d1248):

it says which one is NOT

OpenStudy (haseeb96):

@Ashleyrh1999 guess what i have taught u

OpenStudy (anonymous):

2/5 = 0.4 and all others when divided give 0.36 repeating so 2/5 is the odd one out

OpenStudy (holly00d1248):

i'm goin with 2/5

OpenStudy (jojo4eva):

sorry @Ashleyrh1999 did read the question at first. The answers are 2/5 and 12/33

OpenStudy (anonymous):

There is one answer and is 2/5

OpenStudy (anonymous):

Use a calculator to check

OpenStudy (jojo4eva):

Yup sorry its just 2/5 i meant to say its not 12/33 cuz i had already said it was sorry @familyguymath we make mistakes sometimes

OpenStudy (anonymous):

well, these are all equivalent fractions to 8/22, so now you just need to see which fraction is not in the list and that would be your answer. :) 4/11 = 8/22 = 12/33 = 16/44 = 20/55 = 24/66 = 28/77 = 32/88 = 36/99 = 40/110 = 44/121 = 48/132 = 52/143 = 56/154 = 60/165 = 64/176 = 68/187 = 72/198 = 76/209 = 80/220

OpenStudy (anonymous):

@Ashleyrh1999